# secReferences

### v1

@@ 4.3 Training Objectives
$\mathcal{L}^{RGB/Dep}_{D}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{RGB/Dep}}^{i},{F_{D}^{RGB/Dep}}^{i})\}$ (4)
where $\mathcal{L}^{RGB}_{D}$ and $\mathcal{L}^{Dep}_{D}$ are the losses corresponding to RGB and depth branches respectively.
Secondly, to ensure the effectiveness of Crossmodal Filter, it is crucial to align the mapping features $F_{BP}^{Dep/RGB}$ within these filters to the encoder features. The training losses of RGB and Depth branches are calculated as:
$\mathcal{L}^{RGB/Dep}_{CF}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{RGB/Dep}}^{i},{F_{BP}^{Dep/RGB}}^{i})\}$ (5)
@@
$\mathcal{L}^{RGB/Dep}_{CA}=\mathcal{L}^{RGB/Dep}_{IBP}+\mathcal{L}^{RGB/Dep}_{output}$ (6)
$\mathcal{L}^{RGB/Dep}_{IBP}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{RGB/Dep}}^{i},{F_{IBP}^{Dep/RGB}}^{i})\}$ (7)
$\mathcal{L}^{RGB/Dep}_{output}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{RGB/Dep}}^{i},{F_{CA}^{RGB/Dep}}^{i})\}$ (8)
@@
$\begin{split}\mathcal{L}_{CRD}&=(\mathcal{L}^{RGB}_{D}+\mathcal{L}^{RGB}_{CF}+\mathcal{L}^{RGB}_{CA})\\ &+(\mathcal{L}^{Dep}_{D}+\mathcal{L}^{Dep}_{CF}+\mathcal{L}^{Dep}_{CA})\end{split}$ (9)
@@ 4.4 Anomaly Scoring
$\mathcal{M}^{RGB/Dep}=\sum_{1}^{3}(1-\mathrm{Sim}({F_{E}^{RGB/Dep}}^{i},{F_{CA}^{RGB/Dep}}^{i}))$ (10)
For the fusion of anomaly maps of two branches, we refer to the method in [8]:
  • First, we calculate the means and standard deviations of the values in the anomaly maps using all normal samples from the validation set, denoted as $\mu^{RGB/Dep}$ and $\sigma^{RGB/Dep}$, respectively.
  • Then, during inference, we normalize the anomaly maps for both branches as
    $\mathcal{M}^{RGB/Dep}_{norm}=\frac{\mathcal{M}^{RGB/Dep}-\mu^{RGB/Dep}}{\sigma^{RGB/Dep}}$ (11)
  • Finally, we rescale the maps to the same scale and sum the anomaly maps from both branches to obtain the final anomaly map as $\mathcal{M}=\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$.
And we simply choose $\mathcal{S}=\max(\mathcal{M})$ as the anomaly score.
DepthGAN [7] 42.1 42.2 77.8 69.6 49.4 25.2 28.5 36.2 40.2 63.1 47.4
DepthAE [7] 43.2 15.8 80.8 49.1 84.1 40.6 26.2 21.6 71.6 47.8 48.1
DepthVM [7] 38.8 32.1 19.4 57.0 40.8 28.2 24.4 34.9 26.8 33.1 33.5
BTF [18] 97.6 96.9 97.9 97.3 93.3 88.8 97.5 98.1 95.0 97.1 95.9
M3DM [28] 97.0 97.1 97.9 95.0 94.1 93.2 97.7 97.1 97.1 97.5 96.4
Shape-guided [11] 98.1 97.3 98.2 97.1 96.2 97.8 98.1 98.3 97.4 97.5 97.6
AST [24] 97.0 94.7 98.1 93.9 91.3 90.6 97.9 98.2 88.9 94.0 94.4
MMRD [15] 98.6 99.0 99.1 95.1 99.0 90.1 99.0 99.0 98.7 98.2 97.6
CFM [12] 98.0 96.6 98.2 94.7 95.9 96.7 98.2 98.3 97.6 98.2 97.2
CRD (Ours) 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
Table 2: Pixel-level anomaly localization results PRO (%) on MVTec 3D-AD with the best in bold.
@@ 5.1 Experimental Settings
We conduct experiments on MVTec 3D-AD [7] dataset which is a 3D industrial anomaly detection dataset that includes 10 categories of industrial objects. The training set contains a total of 2,656 normal samples, each consisting of an RGB image and corresponding 3D scan information. The test set includes 1,197 samples, comprising both 948 anomalous samples and 249 normal samples, and provides pixel-level annotations for anomaly detection and localization evaluation. Besides, a validation set is provided by MVTec 3D-AD, containing 294 normal samples.
Evaluation Metrics
@@
The evaluation metrics for anomaly detection include image-level area under the receiver operating characteristic curve (I-AUROC) and average precision (I-AP). For anomaly localization, pixel-level AUROC (P-AUROC) and AP (P-AP) are used. We also employ the per-region overlap (PRO) metric calculated according to the official code [7] for anomaly localization, which treats anomaly regions of different size equally.
Implementation Details
@@
During both training and inference, only the depth information from the 3D data is used. All RGB images and depth maps are scaled to $256\times 256$. The training is performed on a single Nvidia GTX 3090 GPU, and a separate detection model is trained for each category. The network backbone is WideResNet 50. The batch size is 16, and the training lasts for 200 epochs. Each branch is optimized using a separate Adam optimizer with a learning rate of 0.005. In inference, the anomaly maps of both branches are smoothed using a Gaussian filter with $\sigma=4$.
Figure 4: Qualitative results on MVTec-3D AD. From top to bottom: the input images, the corresponding depth maps, the ground truth masks, the output anomaly maps of our proposed CRD.
5.2 Results on MVTec-3D AD
@@
Tab. 1 and Tab. 2 show the quantitative image-level anomaly detection results and pixel-level anomaly localization results on the MVTec 3D-AD dataset. We compare our method, CRD, with several state-of-the-art 3D anomaly detection methods. CRD consistently outperforms all other methods in both I-AUROC and PRO metrics, achieving an average I-AUROC of 96.2% and an average PRO of 97.7%. Notably, CRD demonstrates superior performance in detecting and localizing anomalies across a wide range of categories, including some challenging categories such as Cookie and Peach.
A subset of qualitative results on MVTec-3D AD is provided in Fig. 4, showing that our proposed CRD is able to accurately localize the anomaly regions. For further details on P-AUROC results and more qualitative results please refer to the Supplementary Materials.
@@ 5.3 Ablation Studies
Study on Multi-branch Distillation
Branch I-AUROC P-AUROC P-PRO I-AP P-AP
RGB 86.8 98.9 95.4 95.8 37.7
RGB+Crossmodal 88.9 99.0 95.8 96.5 39.4
Depth 87.1 96.2 88.8 95.2 33.7
Depth+Crossmodal 87.3 97.1 90.1 95.7 34.1
RGB+Depth 93.9 99.3 97.4 98 45.2
RGB+Depth+Crossmodal 96.2 99.4 97.7 98.8 45.5
@@
We conduct an ablation study in Tab. 3, which demonstrates that the multi-branch setup, combining both RGB and Depth modalities, significantly outperforms individual modality branches. The I-AUROC metric of 93.9% for using both branches is higher than 86.8% for RGB branch only and 87.1% for depth branch only, highlighting the complementary benefits of modality fusion.
Study on Crossmodal Assistants
CF CA I-AUROC P-AUROC P-PRO I-AP P-AP
- - 93.9 99.3 97.4 98 45.2
- 94.2 99.3 97.5 98.2 45.8
- 95.6 99.4 97.7 98.6 45.5
96.2 99.4 97.7 98.8 45.5
@@
We perform an ablation study to evaluate the impact of Crossmodal Filter and Crossmodal Amplifier on anomaly detection performance, as shown in Tab. 4. The results reveal that both CF and CA contribute to improved performance, with the combination of both achieving the best results. Specifically, adding CF to the baseline improves I-AUROC from 93.9% to 94.2%, and incorporating CA also boosts the performance, achieving an I-AUROC of 95.6%. The highest performance is achieved when both CF and CA are used together, resulting in an I-AUROC of 96.2%.
Study on Bottlneck Projection in Crossmodal Filter
Downsampling Size I-AUROC P-AUROC P-PRO I-AP P-AP
Identity 93.9 99.3 97.4 98.0 45.8
$16\times 16$ 93.6 99.3 97.4 97.9 45.2
$8\times 8$ 94.2 99.3 97.5 98.2 45.8
$4\times 4$ 93.8 99.3 97.4 98.0 45.3
@@
We investigate the effect of different downsampling sizes on the performance of the Crossmodal Filter, as shown in Tab. 5. When Bottleneck Projection reduces the feature size to $8\times 8$, the model achieves the best performance across multiple metrics. It is observed that excessively large downsampling size (e.g., $4\times 4$) leads to information loss, while too small downsampling size (e.g., $16\times 16$) makes it difficult to filter out abnormal interference effectively. Therefore, we conclude that selecting an appropriate downsampling size is crucial for maintaining a balance between information preservation and interference removal.
Study on Inverted Bottlneck Projection in Crossmodal Amplifier
Channel Expansion I-AUROC P-AUROC P-PRO I-AP P-AP
Identity 95.7 99.3 97.7 98.7 45.1
2×2$\times$2 × 95.6 99.4 97.7 98.6 45.5
4×4$\times$4 × 95.7 99.4 97.7 98.6 45.0
@@
As recorded in Tab. 6, we analyze the impact of different channel expansion sizes on the performance of the Crossmodal Amplifier. The results demonstrate that a 2 times channel expansion strikes the best balance between detection performance and computational efficiency. This configuration achieves a high PRO of 97.7% and a P-AP of 45.5%, while maintaining relatively low computational overhead compared to the 4 times expansion.
Study on Anomaly Scoring
Aggregation I-AUROC P-AUROC P-PRO I-AP P-AP
$\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 96.2 99.4 97.7 98.8 45.5
$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 95.6 99.1 97.6 98.6 44.4
$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 94.3 99.3 97.2 98.2 43.5
@@
Tab. 7 demonstrates that normalizing the anomaly maps from each modality before summing them performs better than directly adding or multiplying the maps. Normalization ensures the anomaly maps from different modalities are on the same scale, making fusion more effective. As a result, normalizing before aggregation improves performance across all metrics.
@@ 6 Conclusion
In this paper, Crossmodal Reverse Distillation is proposed, which offers a novel solution based on Knowledge Distillation for unsupervised multimodal industrial AD task by fully utilizing the information from all modalities. CRD uses a multi-branch design to independently process each modality, enabling more accurate anomaly detection in each branch. Additionally, two crossmodal assistants are introduced: Crossmodal Filter helps the student decoder generate normal features, and Crossmodal Amplifier aims to amplify anomalies of one modality to branches of other modalities. Experimental results on MVTec 3D-AD show that CRD achieves SOTA performance in both anomaly detection and localization.
@@
Appendix A Motivation
Figure 5: Feature diagrams in the anomaly detection process based on Knowledge Distillation. The upper part illustrates previous methods (using RD framework as an example), while the lower part represents the feature transformation in our proposed method CRD. Green and orange indicate two modalities, while blue represents the fused results. In the resulting anomaly maps, darker gray regions indicate higher anomaly scores.
In Sec. 4.1, In Sec 4.1, we introduce the detection approach of previous unsupervised multimodal anomaly detection methods based on Knowledge Distillation. These methods detect and localize anomalies by analyzing the differences between the fused multimodal teacher features and student features. We also identify a key issue in these methods: the anomaly smoothing problem. When features from a normal modality dominate the fusion, abnormal information may be smoothed out, resulting in fused teacher features that fail to effectively reflect the anomalies.
This section provides an illustrative diagram to further explain this problem. In Fig. 5, the upper part outlines the feature transformation process of previous KD-based methods, while the lower part illustrates the feature transformation process in our proposed method. Unlike prior methods, which may miss anomalies when one modality is normal and the other is abnormal, our method ensures sensitivity to anomalies in each modality through dual-branch processing. Additionally, the proposed crossmodal assistants facilitate interaction between modalities, enabling more precise anomaly detection and localization.
Appendix B P-AUROC on on MVTec 3D-AD

### v2

@@ 4.3 Training Objectives
$\mathcal{L}^{2D/3D}_{D}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{D}^{2D/3D}}^{i})\}$ (4)
where $\mathcal{L}^{2D}_{D}$ and $\mathcal{L}^{3D}_{D}$ are the losses corresponding to 2D and 3D branches respectively.
Secondly, to ensure the effectiveness of Crossmodal Filter, it is crucial to align the mapping features $F_{BP}^{3D/2D}$ within these filters to the encoder features. The training losses of 2D and 3D branches are calculated as:
$\mathcal{L}^{2D/3D}_{CF}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{BP}^{3D/2D}}^{i})\}$ (5)
@@
$\mathcal{L}^{2D/3D}_{CA}=\mathcal{L}^{2D/3D}_{IBP}+\mathcal{L}^{2D/3D}_{output}$ (6)
$\mathcal{L}^{2D/3D}_{IBP}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{IBP}^{3D/2D}}^{i})\}$ (7)
$\mathcal{L}^{2D/3D}_{output}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{CA}^{2D/3D}}^{i})\}$ (8)
@@
$\mathcal{L}_{CRD}=(\mathcal{L}^{2D}_{D}+\mathcal{L}^{2D}_{CF}+\mathcal{L}^{2D}_{CA})+(\mathcal{L}^{3D}_{D}+\mathcal{L}^{3D}_{CF}+\mathcal{L}^{3D}_{CA})$ (9)
@@ 4.4 Anomaly Scoring
$\mathcal{M}^{2D/3D}=\sum_{1}^{3}(1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{CA}^{2D/3D}}^{i}))$ (10)
For the fusion of anomaly maps of two branches, we refer to the method in [8]: First, we calculate the means and standard deviations of the values in the anomaly maps using all normal samples from the validation set, denoted as $\mu^{2D/3D}$ and $\sigma^{2D/3D}$, respectively. Then, during inference, we normalize the anomaly maps for both branches as
$\mathcal{M}^{2D/3D}_{norm}=\frac{\mathcal{M}^{2D/3D}-\mu^{2D/3D}}{\sigma^{2D/3D}}$ (11)
Finally, we rescale the maps to the same scale and sum the anomaly maps from both branches to obtain the final anomaly map as $\mathcal{M}=\mathcal{M}^{2D}_{norm}+\mathcal{M}^{3D}_{norm}$. And we simply choose $\mathcal{S}=\max(\mathcal{M})$ as the anomaly score.
M3DM [31] 62.4/90.6 95.8/92.3 95.8/80.3 100/98.3 88.6/85.5 75.8/68.8 94.9/88.0 83.6/90.6 100/96.6 100/95.5 89.7/88.2
AST [26] 57.4/51.4 74.7/83.5 74.7/71.4 88.9/90.5 59.6/58.7 61.7/59.0 81.6/73.6 84.1/76.9 98.7/91.8 98.7/87.8 78.0/74.4
MMRD [15] 85.4/97.5 100/97.0 94.6/94.2 99.8/98.5 90.8/91.7 74.7/68.0 96.6/97.0 98.4/94.1 100/99.0 100/99.2 94.0/93.6
CFM [12] 68.0/94.2 93.1/90.2 95.2/83.1 88.0/96.5 86.5/87.5 78.2/76.2 91.7/79.1 84.0/91.3 99.8/93.9 96.2/94.9 88.1/88.7
CRD (Ours) 86.7/98.0 100/94.1 98.1/96.8 98.1/98.1 87.2/93.6 72.8/77.3 97.1/98.1 93.0/94.4 100/98.9 99.5/98.1 93.3/94.7
Table 2: Anomaly detection and localization results I-AUC/PRO (%) on Eyecandies with the best in bold and the second-best underlined.
@@ 5.1 Experimental Settings
Dataset. We conduct experiments MVTec 3D-AD [7] and Eyecandies [9]. MVTec 3D-AD is a 3D industrial anomaly detection dataset featuring 10 categories of industrial objects. The training set consists of 2,656 normal samples, each including an RGB image and corresponding 3D scan information, while the validation set contains 294 normal samples. The test set comprises 1,197 samples, including 948 anomalous samples and 249 normal samples, with pixel-level annotations provided for evaluation. For our experiments, we use the processed depth maps derived from the 3D scans as the 3D modality. Eyecandies is a synthetic dataset of 10 candy categories in an industrial conveyor scenario, providing RGB images, depth maps, and surface normal maps. It contains 10k training samples, 1k validation samples, and 4k test samples. In our experiments, we select normal maps as the 3D modality and chose images under consistent lighting conditions as the 2D modality.
@@
Evaluation Metrics. The evaluation metrics for anomaly detection include image-level area under the receiver operating characteristic curve (I-AUC) and average precision (I-AP). For anomaly localization, pixel-level AUC (P-AUC) and AP (P-AP) are used. We also employ the per-region overlap (PRO) metric calculated according to the official code [7] for anomaly localization, which treats anomaly regions of different size equally.
@@
Implementation Details. All RGB images, depth maps and normal maps are scaled to $256\times 256$. The training is performed on a single Nvidia GTX 3090 GPU, and a separate detection model is trained for each category. The network backbone is WideResNet-50 [33]. The batch size is 16, and the training lasts for 200 epochs. Each branch is optimized using a separate Adam optimizer with a learning rate of 0.005. In inference, the anomaly maps of both branches are smoothed using a Gaussian filter with $\sigma=4$.
Figure 5: Qualitative ablation results of crossmodal assistants on MVTec 3D-AD.
5.2 Main Results
@@
Results on MVTec 3D-AD. Tab. 1 shows the quantitative image-level anomaly detection results and pixel-level anomaly localization results on the MVTec 3D-AD dataset. We compare our method CRD with several state-of-the-art 3D anomaly detection methods. CRD consistently outperforms all other methods in both I-AUC and PRO metrics, achieving an average I-AUC of 96.2% and an average PRO of 97.7%. Notably, CRD demonstrates superior performance in detecting and localizing anomalies across a wide range of categories, including some challenging categories such as Cookie and Peach. A subset of qualitative results on MVTec 3D-AD is provided in Fig. 1, showing that our proposed CRD accurately localizes the anomaly regions.
Results on Eyecandies. We evaluate our method on the Eyecandies dataset, as shown in Tab. 2. Our method CRD significantly outperforms all non-KD-based multimodal anomaly detection methods. While it slightly lags behind MMRD [15] which uses both depth maps and surface normal maps as 3D modalities in anomaly detection metrics I-AUC, it consistently achieves superior performance 94.7% in anomaly localization metrics PRO.
Method MVTec 3D-AD (%) Eyecandies (%) Avg. (%) Frame Rate (fps)
I-AUC P-AUC PRO I-AUC P-AUC PRO
AST [26] 93.7 97.6 94.4 78.0 90.2 74.4 88.1 13.9
MMRD* [15] 95.0 99.2 97.6 94.0 98.3 93.6 96.3 -
CFM [12] 96.0 99.3 97.2 88.1 97.4 88.7 94.5 12.0
CRD (Ours) 96.2 99.4 97.7 93.3 98.5 94.7 96.6 21.7
Table 3: Comparison of feature learning-based methods on MVTec 3D-AD and Eyecandies datasets under the same hardware setup. * indicates that MMRD has no reproducible code.
Comparison with Feature Learning-based Methods. We further compare CRD with feature learning-based methods including AST[26], MMRD[15], and CFM [12] on MVTec 3D-AD and Eyecandies. As in Table 3, CRD achieves the best average performance 96.6% and the highest frame rate 21.7 fps. Despite the multi-branch design, CRD remains lightweight due to its CNN-based architecture, outperforming other feature learning-based methods in efficiency.
@@ 5.3 Ablation Studies
Branch I-AUC P-AUC P-PRO I-AP P-AP Avg.
2D 86.8 98.9 95.4 95.8 37.7 82.9
2D+Crossmodal 88.9 99.0 95.8 96.5 39.4 83.9
3D 87.1 96.2 88.8 95.2 33.7 80.2
3D+Crossmodal 87.3 97.1 90.1 95.7 34.1 80.9
2D+3D 93.9 99.3 97.4 98.0 45.2 86.8
2D+3D+Crossmodal 96.2 99.4 97.7 98.8 45.5 87.5
@@
CF CA I-AUC P-AUC P-PRO I-AP P-AP Avg.
- - 93.9 99.3 97.4 98.0 45.2 86.8
- 94.2 99.3 97.5 98.2 45.8 87.0
- 95.6 99.4 97.7 98.6 45.5 87.4
96.2 99.4 97.7 98.8 45.5 87.5
@@
Study on Multi-branch Distillation. We conduct an ablation study in Tab. 4, which demonstrates that the multi-branch setup, combining both 2D and 3D modalities, significantly outperforms individual modality branches. The I-AUC metric of 93.9% for using both branches is higher than 86.8% for 2D branch only and 87.1% for 3D branch only, highlighting the complementary benefits of modality fusion.
Study on Crossmodal Assistants. We perform an ablation study to evaluate the impact of crossmodal assistants on anomaly detection performance, as shown in Tab. 5. The results reveal that both CF and CA contribute to improved performance, with the combination of both achieving the best results. Specifically, adding CF to the baseline improves the I-AUC from 93.9% to 94.2%, and incorporating CA also boosts the performance, achieving an I-AUC of 95.6%. The highest performance is achieved when both CF and CA are used together, resulting in an I-AUC of 96.2%.
We also provide visualizations of anomaly maps before and after incorporating crossmodal assistants, as in Fig. 5. Without crossmodal assistants, anomalies that are less prominent in a specific modality goes undetected in the corresponding branch. However, after introducing crossmodal assistants, this issue is significantly mitigated, resulting in a noticeable reduction in the anomaly miss rate across all modality branches.
Downsampling Size I-AUC P-AUC P-PRO I-AP P-AP Avg.
Identity 93.9 99.3 97.4 98.0 45.8 86.9
$16\times 16$ 93.6 99.3 97.4 97.9 45.2 86.7
$8\times 8$ 94.2 99.3 97.5 98.2 45.8 87.0
$4\times 4$ 93.8 99.3 97.4 98.0 45.3 86.8
@@
Study on Bottlneck Projection in Crossmodal Filter. We investigate the effect of different downsampling sizes on the performance of the Crossmodal Filter, as shown in Tab. 6. When Bottleneck Projection reduces the feature size to $8\times 8$, the model achieves the best performance across multiple metrics. It is observed that excessively large downsampling size (e.g., $4\times 4$) leads to information loss, while small downsampling size (e.g., $16\times 16$) makes it difficult to filter out abnormal interference effectively. Therefore, we conclude that selecting an appropriate downsampling size is crucial for maintaining a balance between information preservation and interference removal.
Channel Expansion I-AUC P-AUC P-PRO I-AP P-AP Avg.
Identity 95.7 99.3 97.7 98.7 45.1 87.3
2×2$\times$2 × 95.6 99.4 97.7 98.6 45.5 87.4
4×4$\times$4 × 95.7 99.4 97.7 98.6 45.0 87.3
@@
Study on Inverted Bottlneck Projection in Crossmodal Amplifier. As recorded in Tab. 7, we analyze the impact of different channel expansion sizes on the performance of the Crossmodal Amplifier. The results demonstrate that a 2 times channel expansion strikes the best balance between detection performance and computational efficiency. This configuration achieves a high PRO of 97.7% and a P-AP of 45.5%, while maintaining relatively low computational overhead compared to the 4 times expansion.
Aggregation I-AUC P-AUC P-PRO I-AP P-AP Avg.
$\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 96.2 99.4 97.7 98.8 45.5 87.5
$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 95.6 99.1 97.6 98.6 44.4 87.1
$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 94.3 99.3 97.2 98.2 43.5 86.5
@@
Study on Anomaly Scoring. Tab. 8 demonstrates that normalizing the anomaly maps from each modality before summing them performs better than directly adding or multiplying the maps. Normalization ensures that anomaly maps from different modalities are on the same scale, making fusion more effective.
Figure 6: Visualization of the features output by CF and CA.
Visualization of Mapped Features. We visualizes the output features of CF and CA using the visualization decoders [32] in Fig. 6, which intuitively demonstrates the roles of CA and CF. Regardless of whether the initial modality is normal or anomalous, CF consistently produces normal outputs, enabling the student decoder of the other modality to generate anomaly-free results. When the initial modality is anomalous, CA introduces discrepancies with the other modality, amplifying anomalies for enhanced detection.
@@ 6 Conclusion
In this paper, Crossmodal Reverse Distillation is proposed, which offers a novel solution based on Knowledge Distillation for unsupervised multimodal AD task. CRD uses a multi-branch design to independently process each modality, enabling more accurate anomaly detection in each branch. Additionally, two crossmodal assistants are introduced: Crossmodal Filter helps the student decoder generate normal features, and Crossmodal Amplifier aims to amplify anomalies of one modality to branches of other modalities. Experimental results show that CRD achieves SOTA performance in both anomaly detection and localization.
@@
Overview
This supplementary material provides extended analyses and visualizations to complement the main paper:
  • Detailed P-AUC results for anomaly localization on MVTec 3D-AD and Eyecandies datasets. (Appendix A)
  • Training accuracy curves to demonstrate the model’s convergence over training. (Appendix B)
  • Detailed results of ablation studies on MVTec 3D-AD dataset. (Appendix C)
  • Additional qualitative visualizations on MVTec 3D-AD and Eyecandies datasets. (Appendix D)
Appendix A P-AUC Metric
A.1 P-AUC on MVTec 3D-AD